How much Volume of 0.1 M  of KMnO4 required to react completely with 100ml of 0.1 FeSO4 solution in acidic medium ?

Grade:11

1 Answers

Sahani Kumar
5 years ago
In acidic medium M7+  get reduced to M2+
                       M7+  +5e- --------> M2+        so n1=5
 
Given:M1=0.1    M2=0.1
            V1=?        V2=100
 
In acidic medium Fe2+ get oxidised to Fe3+
                       Fe2+ ------------->Fe3+  + 1e-      so n2= 1
Now,  we know that 
                      N1 * V1=N2 * V2
             n1 * M1 * V1=n2 * M2 * V2
               5 * 0.1 * V1=1 * 0.1 * 100
                              V1=(1*0.1*100)/(5*0.1)
                              V1=20ml